# A cutting-surface consensus approach for distributed robust optimization of multi-agent systems

### v2

@@ 4.1 Distributed Cutting-Surface Consensus Approach
1:  Solve ($\rm{ADRCP}^{k}$) by the DPG algorithm equipped with distributed termination algorithm proposed in Section 3.
2:  if solvable then
3:     $\bullet$ Set $x^{k}_{i}$ equal to the ($\rm{ADRCP}^{k}$) approximate optimal solution point for agent $i$.
4:     $\bullet$ Solve (${\rm LLP}_{i}$) to global optimality for agent $i$, and obtain $g_{i}^{\max}(x^{k}_{i})$ and $y^{\max}_{i}$.
$g^{\max}_{i}(x^{k}_{i})\!=\!g_{i}(x^{k}_{i},y^{\max}_{i})\!=\!\mathop{\max}\limits_{{y_{i}}\in Y_{i}}g_{i}(x^{k}_{i},y_{i})$ (${\rm LLP}_{i}$)
5:     if $g_{i}^{\max}(x^{k}_{i})\leq 0$ then
6:        $\ast$ Update the local feasible point of agent $i$ for the (${\rm DRCP}$) problem: $z_{i}^{k}=x^{k}_{i}$.
7:        $\ast$ Execute Algorithm 2 to check whether the following stopping criterion is satisfied or not: for any two agents $i\neq j$, and $i,j\in\mathcal{V}$
$\|z_{i}^{k}-z_{j}^{k}\|\leq\epsilon_{4}$ (9a)
$\|z_{i}^{k}-z_{i}^{k-1}\|\leq\epsilon_{5}$ (9b)
$|f_{i}(z_{i}^{k})-f_{i}(z_{i}^{k-1})|\leq\epsilon_{6}$ (9c)
8:        if stopping criterion (9) is satisfied then
9:           Output $z_{i}^{k}$ for all $i\in\mathcal{V}$ and Terminate.
10:        end if
11:        $\ast$ Optimality cuts: set $Y^{k+1}_{i}\!\leftarrow\!Y^{k}_{i}$, $\epsilon^{k+1}_{i}\!\leftarrow\!\epsilon^{k}_{i}/r$
@@
13:        $\ast$ Set the point $z_{i}^{k}=z_{i}^{k-1}$
14:        $\ast$ Feasibility cuts: set $Y^{k+1}_{i}\!\!\leftarrow\!Y^{k}_{i}\cup\left\{y^{\max}_{i}\right\}\!,\ \!\!\epsilon^{k+1}_{i}\!\!\leftarrow\!\epsilon^{k}_{i}$
15:     end if
16:  else
17:     $\bullet$ Solvability cuts: set $Y^{k+1}_{i}\!\leftarrow\!Y^{k}_{i},\epsilon^{k+1}_{i}\!\leftarrow\!\epsilon^{k}_{i}/r,\forall i\!\in\!\mathcal{V}$
18:  end if
19:  Set $k\leftarrow k+1$
@@ 4.1.1 The first selection structure
The ($\rm{ADRCP}^{k}$) is neither a relaxation nor a restriction of the (${\rm DRCP}$) in general. In the initialization process, we choose any initial finite or empty set $Y_{i}^{0}\subset Y_{i}$ and any initial restriction parameter $\epsilon^{0}_{i}>0$ for each agent. Since there are relatively large restriction parameters $\epsilon^{k}_{i}$ in the first few iterations of Algorithm 1, the global feasible set $\mathcal{F}^{k}$ of the ($\rm{ADRCP}^{k}$) may be an empty set, in which case, specifically, there exist two situations for the ($\rm{ADRCP}^{k}$): one is that the local feasible set of a certain agent is empty, and the other is that the local feasible sets of all agents are non-empty while their intersection is empty. We use the DRG algorithm to solve the ($\rm{ADRCP}^{k}$). For the first situation, each agent is forced to terminate and cannot obtain a solution. For the second situation, all agents update their states infinitely since the consensus stopping condition (4a) fails to be satisfied.
In Section 3.2, we combine the DPG algorithm with the finite-time termination method to enable each agent to find an approximate optimal and consensus solution of the ($\rm{ADRCP}^{k}$) within finite time slots under Assumption 3. In order to terminate the DPG algorithm for any combinations of $Y^{k}_{i}$ and $\epsilon^{k}_{i}$ in finite time slots without Assumption 3, we set a maximum number of time slots $T$ on the basis of the above algorithm. When the DPG algorithm cannot yield a solution or fail to reach the stopping criterion (4) within the time slot $T$, we can determine that the ($\rm{ADRCP}^{k}$) is unsolvable. Therefore, we relax the ($\rm{ADRCP}^{k}$) by proportionally decreasing the restriction parameters $\epsilon_{i}^{k}$ for all agents (Solvability cuts). If the DPG algorithm terminates by satisfying the stopping criterion (4) within the time slot $T$, then we obtain the approximate optimal solution point for the ($\rm{ADRCP}^{k}$) and update $x_{i}^{k}$.
@@ 4.1.2 The second selection structure
When the ($\rm{ADRCP}^{k}$) is solvable, the obtained solution $x^{k}_{i}$ of agent $i$ has two cases. One is that $x^{k}_{i}$ is in its local feasible region of the (${\rm DRCP}$), i.e., $x^{k}_{i}\!\in\!\mathcal{F}_{i}$, the other is that $x^{k}_{i}$ is not in agent $i$’s feasible region of the (${\rm DRCP}$), i.e., $x^{k}_{i}\!\notin\!\mathcal{F}_{i}$. To judge whether $x^{k}_{i}$ belongs to $\mathcal{F}_{i}$, we introduce the lower level program (${\rm LLP}_{i}$).
If $g^{max}_{i}(x^{k}_{i})\leq 0$, it indicates that $x^{k}_{i}\in\mathcal{F}_{i}$. Then, we assign the value of $x_{i}^{k}$ to the (${\rm DRCP}$) optimal solution candidate $z_{i}^{k}$, and reduce $\epsilon^{k}_{i}$ proportionally to relax the ($\rm{ADRCP}^{k}$) problem which is aimed to obtain a solution that is closer to the optimal solution point of the (${\rm DRCP}$) in later iterations (Optimality cuts). If $x_{i}^{k}\notin\mathcal{F}_{i}$, the obtained $x_{i}^{k}$ has no practical significance and needs to be discarded. At this time, we expand the point $y^{\max}_{i}=\arg{\max}_{{y_{i}}\in Y_{i}}g_{i}(x^{k}_{i},y_{i})$ to $Y_{i}^{k}$ to tighten the ($\rm{ADRCP}^{k}$), aiming to make $x^{k}_{i}$ locally feasible for the (${\rm DRCP}$) in later iterations (Feasibility cuts). This process is illustrated by a numerical example [^1] Example: $\mathop{\min}_{x\in\mathcal{F}_{i}}-x(2)\quad\ {\rm subject\ to:}\ g_{i}(x,y_{i})\leq 0,\ \forall y_{i}\in Y_{i},$
where $x=[x(1),x(2)]^{\top}$, $\mathcal{F}_{i}=[0,2]\times[0,1]$, $Y_{i}=[0,2]$, and $g_{i}(x,y_{i})=x(2)+[x(1)^{2}\!-\!2x(1)]e^{-{x(1)}^{2}+{y_{i}}^{2}-2x(1)y_{i}}$.
, as shown in Figure 2.
Figure 2: Grahpic illustration of the process in the second selection structure. The region enclosed by the red dashed line represents the (DRCP) feasible region. The three regions enclosed by the solid lines correspond to the (${\rm ADRCP}^{k}$) feasible regions formed by three consecutive iterations of Algorithm 1. $x^{k}_{i}$ is the approximate optimal solution of the (${\rm ADRCP}^{k}$). $x^{1}_{i}$ is feasible to the (DRCP), thus an optimality cut is performed to reduce the restriction parameter $\epsilon^{2}_{i}=\epsilon^{1}_{i}/r=0.1$. $x^{2}_{i}$ is infeasible to the (DRCP), thus a feasibility cut is performed to expand the finite set $Y^{3}_{i}=Y^{2}_{i}\cup\{y^{\max}_{i}\}=\{1,0.3\}$. At the third iteration of Algorithm 1, the resulting solution $x^{3}_{i}$ of the (${\rm ADRCP}^{k}$) returns to feasible for the (DRCP).
@@
where $x^{k}_{i}$ is fixed, $y_{i}\in Y_{i}$, $\dot{\chi}(y_{i})$ represents the derivative of the function $\chi(y_{i})$, and $\dot{g}_{i}(x^{k}_{i},y_{i})$ represents the derivative of the function $g_{i}(x^{k}_{i},y_{i})$ with respect to $y_{i}$. Set the compact set $Y_{i}=\left\{y_{i}|y_{0}\leq y_{i}\leq y_{f}\right\}$ and $\chi(y_{0})=g_{i}(x^{k}_{i},y_{0})$. Then, adopting an adaptive step-size solver with rigorous state-event location, we obtain the maximum constraint violation ${\max}_{y_{i}\in Y_{i}}\ g_{i}(x^{k}_{i},y_{i})=\chi(y_{f})$. Moreover, $y^{\max}_{i}$ is obtained from the last event triggered in the integration process.
4.1.3 The third selection structure
To ensure finite-time termination of Algorithm 1, we adapt the distributed finite-time termination algorithm in Section 3.2 to verify the stopping criterion (9), where (9a) refers to the consensus stopping condition, while (9b) and (9c) denote the zeroth-order stopping conditions based on the candidate approximate solution $z_{i}^{k}$ and function values $f_{i}(z_{i}^{k})$ respectively. The pseudocode is shown in Algorithm 2.
0:  time slot $t=0$; parameter $e^{4}_{i}(t)=0$; $e^{5}_{i}(t)=0$; $e^{6}_{i}(t)=0$; $h^{2}_{i}(t)=0$.
0:  optimal solution candidate $z_{i}^{k}$ for all agents; parameter $S$ of the UJSC network; the diameter $D$ of $\mathcal{G}(t:t+S)$; certain tolerances $\epsilon_{4}$, $\epsilon_{5}$, $\epsilon_{6}$.
@@ 4.1.3 The third selection structure
1:  Local information exchange: send $z_{i}^{k}$, $z_{i}^{k-1}$, $e^{4}_{i}(t)$, $e^{5}_{i}(t)$, $e^{6}_{i}(t)$ and $h^{2}_{i}(t)$ to the out-neighbors of agent $i$ between the time slot $t$ and $t+1$.
@@
$e^{4}_{i}(t\!+\!1)\!=\!\begin{cases}e^{4}_{i}(t)\!+\!1,\quad\|z_{i}^{k}-z_{j}^{k}\|\!\leq\!\epsilon_{4}\ \\ \qquad\qquad\quad\ \forall j\in{N_{i}^{in}(t)}\cup\left\{i\right\},\\ 0,\quad\ \qquad\quad otherwise,\end{cases}$ (12)
$e^{5}_{i}(t\!+\!1)\!=\!\begin{cases}e^{5}_{i}(t)\!+\!1,\quad\|z_{j}^{k}-z_{j}^{k-1}\|\!\leq\!\epsilon_{5}\ \\ \qquad\qquad\quad\ \forall j\in{N_{i}^{in}(t)}\cup\left\{i\right\},\\ 0,\qquad\ \quad\quad otherwise,\end{cases}$ (13)
$e^{6}_{i}(t\!+\!1)\!=\!\begin{cases}e^{6}_{i}(t)\!+\!1,\!\quad|f_{j}(z_{j}^{k})\!-\!f_{j}(z_{j}^{k-1})|\!\leq\!\epsilon_{6}\\ \qquad\qquad\quad\ \forall j\in{N_{i}^{in}(t)}\cup\left\{i\right\},\\ 0,\qquad\ \quad\quad otherwise.\end{cases}$ (14)
@@
6:  a predefined stopping rule (e.g. the time slot number that is greater than $SD+1$) is satisfied.
@@
Consider Algorithm 2 under Assumption 2. If at the $(SD+1)$-th time slot, there exists an agent $i\in\mathcal{V}$ that satisfies $h^{2}_{i}(t)\geq SD+1$, then all agents meet the stopping criterion (9).
Proof 4.8.
According to Proposition 1, the above result is straightforward.
Algorithm 2 provides Algorithm 1 with termination detection. If there exists an agent satisfying $g_{i}^{\max}(x^{k}_{i})\leq 0$ in Step 5 of Algorithm 1, Algorithm 2 will start to execute. When the network satisfies the stopping criterion (9), all agents spontaneously issue an exit command to terminate Algorithm 1. Besides, an approximate optimal solution $z_{i}^{k}$ of the (${\rm DRCP}$) is output for each agent.
Different from the distributed finite-time termination algorithm introduced in Section 3.2, Algorithm 2 serves primarily to judge the satisfaction of the stopping criterion (9) of the candidate approximate optimal solutions $z_{i}^{k}$ for all agents at the $k$-th iteration, where the value of $z_{i}^{k}$ does not change with respect to the time slots $t$, but it is affected by the iterations $k$. Besides, Algorithm 2 requires only a fixed number of time slots ($SD+1$) to determine whether Algorithm 1 needs to be terminated or not.
@@ 4.2 Convergence Analysis
This subsection presents the main result of Algorithm 1. At first, to ensure convergence of Algorithm 1, we make the following assumption.
@@
$F(\beta x_{1}+(1-\beta)x_{2})<\beta F(x_{1})+(1-\beta)F(x_{2})$
For the case that $F(x)$ is a convex function, we can adopt a tie-break rule in solving the ($\rm{ADRCP}^{k}$) to ensure the uniqueness of the solution.
Proposition 4.9 (Convergence).
[…]
Proof 4.10.
In the loop of Algorithm 1, $\epsilon_{i}^{k}$ is a positive parameter that leads to restriction, while $Y_{i}^{k}$ is a set consisting of a finite number of points obtained from the compact set $Y_{i}$ that results in relaxation. Therefore, the (${\rm ADRCP^{k}}$) is neither a relaxation nor a restriction of the (${\rm DRCP}$) in general. It is difficult for us to directly analyze the relationship between the solutions of the two problems (${\rm ADRCP^{k}}$) and (${\rm DRCP}$). To prove this proposition, we first make the following settings:
[…]
where $\epsilon^{k}_{i}$ and $Y_{i}^{k}$ are the parameters corresponding to the (${\rm ADRCP^{k}}$) at the $k$-th iteration of Algorithm 1. Taking $F(x)$ as the objective function and each of the above four sets as the constraint, we obtain the corresponding optimal solutions $x^{*}$, $x^{k,*}$, $x^{k,a}$, $x^{k,b}$ by solving them in turn. Since the relation $Y^{k}_{i}\subset Y^{k+1}_{i}$ holds for any $k\geq 0$ and $i\in\mathcal{V}$, $Y_{i}$ and $\mathcal{B}^{k}$ are compact, and $g_{i}$ are continuous on $X\times Y_{i}$ for any $i\in\mathcal{V}$, by Lemma 2.2 in [32], it holds that
@@
Since the positive sequence $\left\{\epsilon^{k}_{i}\right\}_{k}$ is nonincreasing for $i\in\mathcal{V}$, it holds that for any $k\geq 0$
$\mathcal{A}^{k}\subset\mathcal{A}^{k+1}\subset\cdots\subset\mathcal{A}^{\infty},$
In addition, since $\epsilon^{k}_{i}\rightarrow 0$ as $k\rightarrow\infty$ for all agents, it follows that $\mathcal{A}^{\infty}=\mathcal{F}$. Based on the this result, we obtain that

### v3

@@ 3.3 Distributed Cutting-Surface Consensus Approach for Solving The DRCP
1: Compute $x^{k}_{i}$ by solving ($\rm{ADRCP}^{\it k}$) using Algorithm 1.
2:if $x^{k}_{i}\in\emptyset$ then
2:
3:  $\bullet$ $\epsilon^{k+1}_{i}\leftarrow\epsilon^{k}_{i}/r$
4:  $\bullet$ $Y^{k+1}_{i}\leftarrow Y^{k}_{i}$
5:else
6:  $\bullet$ $y^{\max,k}_{i}\in\arg\mathop{\max}\limits_{{y_{i}}\in Y_{i}}g_{i}(x^{k}_{i},y_{i})$       $({\rm LLP}_{i})$% Some effective numerical methods for solving $({\rm LLP}_{i})$ are discussed in Remark 3.6.
7:  $\bullet$ $g^{\max,k}_{i}=g_{i}(x^{k}_{i},y^{\max,k}_{i})$
8:  if $g_{i}^{\max,k}>0$ then
8:
9:   $\ast$ $\epsilon^{k+1}_{i}\leftarrow\epsilon^{k}_{i}$
10:   $\ast$ $Y^{k+1}_{i}\leftarrow Y^{k}_{i}\cup\{y^{\max,k}_{i}\}$
10:
11:   $\ast$ $z_{i}^{k+1}\leftarrow z_{i}^{k}$
@@
13:   $\ast$ $z_{i}^{k+1}\leftarrow x^{k}_{i}$.
14:   $\ast$ Check whether stopping criterion (10) is met by using Algorithm 3.
15:   if not then
15:
16:    $\star$ $\epsilon^{k+1}_{i}\leftarrow\epsilon^{k}_{i}/r$
17:    $\star$ $Y^{k+1}_{i}\leftarrow Y^{k}_{i}$
17:
18:   else
19:    $\star$ return $z_{i}^{k+1}$ and Terminate.
20:   end if
21:  end if
22:end if
23: Set $k\leftarrow k+1$
@@ 3.3.1 The first selection structure
The ($\rm{ADRCP}^{\it k}$) is neither a relaxation nor a restriction of the ($\rm DRCP$}) in general. In the input step, we choose any finite or empty set $Y_{i}^{0}\subset Y_{i}$ and any initial restriction parameter $\epsilon^{0}_{i}>0$ for each agent. Since there are relatively large restriction parameters $\epsilon^{k}_{i}$ in the first few iterations of Algorithm 2, the global feasible set $\mathcal{F}^{k}$ of the ($\rm{ADRCP}^{\it k}$) may be an empty set, in which case, specifically, there exist two situations for the ($\rm{ADRCP}^{\it k}$): one is that the local feasible set of a certain agent is empty, and the other is that the local feasible sets of all agents are non-empty while their intersection is empty. We use Algorithm 1 to solve the ($\rm{ADRCP}^{\it k}$). For the first situation, each agent is forced to terminate and cannot obtain a solution. For the second situation, all agents update their local estimates infinitely since the consensus stopping condition (4a) fails to be satisfied.
In order to terminate the DPG algorithm for any combinations of $Y^{k}_{i}$ and $\epsilon^{k}_{i}$ in finite time without Assumption 3, we set maximum iteration time $T$ on the basis of Algorithm 1. When the DPG algorithm cannot yield a solution or fail to reach the stopping criterion (4) within $T$, we can determine that the ($\rm{ADRCP}^{\it k}$) is infeasible. Therefore, we relax the ($\rm{ADRCP}^{\it k}$) by proportionally decreasing the restriction parameters $\epsilon_{i}^{k}$ for all agents. If the DPG algorithm terminates by satisfying the stopping criterion (4) within $T$, then we obtain the approximate optimal solution of the ($\rm{ADRCP}^{\it k}$) and update $x_{i}^{k}$.
@@ 3.3.2 The second selection structure
When the ($\rm{ADRCP}^{\it k}$) is feasible, the obtained solution $x^{k}_{i}$ of agent $i$ has two cases. One is that $x^{k}_{i}$ is in its local feasible region of the ($\rm DRCP$}), i.e., $x^{k}_{i}\!\in\!\mathcal{F}_{i}$, the other is that $x^{k}_{i}$ is not in agent $i$’s feasible region of the ($\rm DRCP$}), i.e., $x^{k}_{i}\!\notin\!\mathcal{F}_{i}$. To judge whether $x^{k}_{i}$ belongs to $\mathcal{F}_{i}$, we introduce a lower level program $({\rm LLP}_{i})$.
If $g^{\max,k}_{i}\leq 0$, it indicates that $x^{k}_{i}\in\mathcal{F}_{i}$. Then, we assign the value of $x_{i}^{k}$ to the ($\rm DRCP$}) optimal solution candidate $z_{i}^{k+1}$, and reduce $\epsilon^{k}_{i}$ proportionally to relax the ($\rm{ADRCP}^{\it k}$) problem which is aimed to obtain a solution that is closer to the optimal solution point of the ($\rm DRCP$}) in later iterations (Optimality cuts). If $x_{i}^{k}\notin\mathcal{F}_{i}$, the obtained $x_{i}^{k}$ has no practical significance and needs to be discarded. At this time, we add the point $y^{\max,k}_{i}$ to $Y_{i}^{k}$ to tighten the ($\rm{ADRCP}^{\it k}$), aiming to make $x^{k}_{i}$ locally feasible to the ($\rm DRCP$}) in later iterations (Feasibility cuts).
@@
Given that $x_{i}^{k}$ is fixed for each agent $i\in\mathcal{V}$, the $({\rm LLP}_{i})$ is a standard optimization problem with a set constraint $y_{i}\in Y_{i}$, where $y_{i}$ is the decision variable. In order to ensure the local feasibility of the solutions generated by Algorithm 2, each agent needs to solve its $({\rm LLP}_{i})$ at global optimality by local computation. [27, 44] used convex numerical methods to locate the maximum violation for the convex $({\rm LLP}_{i})$. When the $({\rm LLP}_{i})$ is nonconvex, [45] presents an effective numerical method under the assumption that $g_{i}$ is continuously differentiable. Specifically, we first introduce an additional function $\chi(y_{i})$:
where $x^{k}_{i}$ is fixed, $y_{i}\in Y_{i}$, $\dot{\chi}(y_{i})$ represents the derivative of the function $\chi(y_{i})$, and $\dot{g}_{i}(x^{k}_{i},y_{i})$ represents the derivative of the function $g_{i}(x^{k}_{i},y_{i})$ with respect to $y_{i}$. Set the compact set $Y_{i}=\left\{y_{i}|y_{0}\leq y_{i}\leq y_{f}\right\}$ and $\chi(y_{0})=g_{i}(x^{k}_{i},y_{0})$. Then, adopting an adaptive step-size solver with rigorous state-event location, we obtain the maximum constraint violation ${\max}_{y_{i}\in Y_{i}}\ g_{i}(x^{k}_{i},y_{i})=\chi(y_{f})$. Moreover, $y^{\max,k}_{i}$ is obtained from the last event triggered in the integration process.
3.3.3 The third selection structure
To ensure finite-time termination of Algorithm 2, we adapt the distributed finite-time termination algorithm in Section 3.2 to verify the following stopping criterion: for any two agents $i\neq j$, and $i,j\in\mathcal{V}$
$\|z_{i}^{k+1}-z_{j}^{k+1}\|\leq\epsilon_{4}$ (10a)
$\|z_{i}^{k+1}-z_{i}^{k}\|\leq\epsilon_{5}$ (10b)
$|f_{i}(z_{i}^{k+1})-f_{i}(z_{i}^{k})|\leq\epsilon_{6}$ (10c)
where (10a) refers to the consensus stopping condition, while (10b) and (10c) denote the zeroth-order stopping conditions based on the optimal solution candidates $z_{i}^{k}$ and function values $f_{i}(z_{i}^{k})$ respectively. The pseudocode is shown in Algorithm 3.
$0h^{2}_{i}(t)=$0, $e^{4}_{i}(t)=0$, $e^{5}_{i}(t)=0$, $e^{6}_{i}(t)=0$, $\forall i\in\mathcal{V}$; $t=0$.
@@ 3.3.3 The third selection structure
1:Local information exchange: send the values $z_{i}^{k}$, $z_{i}^{k+1}$, $f_{i}(z_{i}^{k})$, $f_{i}(z_{i}^{k+1})$, $e^{4}_{i}(t)$, $e^{5}_{i}(t)$, $e^{6}_{i}(t)$ and $h^{2}_{i}(t)$ to the out-neighbors of agent $i$
@@
$e^{4}_{i}(t\!+\!1)\!=\!\begin{cases}e^{4}_{i}(t)\!+\!1,\quad\|z_{i}^{k+1}-z_{j}^{k+1}\|\!\leq\!\epsilon_{4}\ \\ \qquad\qquad\quad\ \forall j\in{N_{i}^{in}(t)}\cup\left\{i\right\},\\ 0,\quad\ \qquad\quad otherwise,\end{cases}$ (12)
$e^{5}_{i}(t\!+\!1)\!=\!\begin{cases}e^{5}_{i}(t)\!+\!1,\quad\|z_{j}^{k+1}-z_{j}^{k}\|\!\leq\!\epsilon_{5}\ \\ \qquad\qquad\quad\ \forall j\in{N_{i}^{in}(t)}\cup\left\{i\right\},\\ 0,\qquad\ \quad\quad otherwise,\end{cases}$ (13)
$e^{6}_{i}(t\!+\!1)\!=\!\begin{cases}e^{6}_{i}(t)\!+\!1,\ |f_{j}(\!z_{j}^{k+1}\!)\!-\!f_{j}(z_{j}^{k})|\!\leq\!\epsilon_{6}\\ \qquad\qquad\ \ \forall j\in{N_{i}^{in}(t)}\cup\left\{i\right\},\\ 0,\qquad\ \quad\ otherwise.\end{cases}$ (14)
@@
6: a predefined stopping rule (e.g. $t>SD+1$) is satisfied.
@@
Consider Algorithm 3 under Assumption 2. If there exists an agent $i\in\mathcal{V}$ that satisfies $h^{2}_{i}(t)\geq SD+1$ at time $SD+1$, then the stopping criterion (10) is satisfied for all agents of the multi-agent system.
Proof 3.8.
Similar to the proof idea of Proposition 1, the above result can be easily derived.
Algorithm 3 provides Algorithm 2 with termination detection. If there exists an agent satisfying $g_{i}^{\max,k}\leq 0$ in Step 13 of Algorithm 2, Algorithm 3 will start to execute. When the network satisfies the stopping criterion (10), all agents spontaneously issue an exit command to terminate Algorithm 2. Additionally, an approximate optimal solution $z_{i}^{k+1}$ of the ($\rm DRCP$}) is output for each agent.
Different from the distributed finite-time termination algorithm introduced in Section 3.2, Algorithm 3 serves primarily to judge the satisfaction of the stopping criterion (10) of the optimal solution candidates $z_{i}^{k}$ for all agents at the $k$-th iteration, where the value of $z_{i}^{k}$ does not change with respect to time $t$, but it is affected by the iteration $k$. In addition, Algorithm 3 requires only a fixed number of time ($SD+1$) to determine whether Algorithm 2 needs to be terminated.
@@ 3.4 Convergence Analysis
This subsection presents the main result of Algorithm 2. Firstly, to ensure finite-time termination of Algorithm 2, we make the following assumption.
@@
The global objective function $F(x)$ of the ($\rm DRCP$}) is strictly convex, i.e., for any two point $x_{1}\neq x_{2}\in X$, $0<\beta<1$, there is
$F(\beta x_{1}+(1-\beta)x_{2})<\beta F(x_{1})+(1-\beta)F(x_{2}).$
Note that for the case that $F(x)$ is a convex function, we can adopt a tie-break rule in solving the ($\rm{ADRCP}^{\it k}$) to ensure the uniqueness of the solution.
Proposition 3.9 (Finite-time Termination).
For each agent $i\in\mathcal{V}$, take any initial restriction parameter $\epsilon_{i}^{0}>0$, any initial finite or empty set $Y^{0}_{i}\subset Y_{i}$, and any reduction parameter $r>1$. Under Assumptions 1, 2, 46, the distributed cutting-surface consensus algorithm terminates in a finite number of iterations and obtain solutions for all agents that satisfy consensus and zeroth-order stopping conditions to certain tolerances.
Proof 3.10.
To prove the finite-time termination of the algorithm, we only need to exclude three possible infinite cases (Case I–III), which are presented in Figure 1.
1) Exclusion of infinite Case I: By Assumption 1, there exists at least one slater point $\hat{x}$ for the ($\rm DRCP$}), we have
where $0<\epsilon_{i}^{s}\leq\min_{y_{i}\in Y_{i}}{-g_{i}(\hat{x},y_{i})},\ \forall i\in\mathcal{V}$. Assume that the (${\rm ADRCP}^{k}$) is infeasible at the $k$-th iteration, i.e., $\mathcal{F}^{k}=\emptyset$. Then, after at most $K_{1}=\max_{i\in\mathcal{V}}\lceil\log_{r}\epsilon^{k}_{i}/\epsilon^{s}_{i}\rceil$ updates of restriction parameters $\epsilon^{k}_{i}$, it follows that $0<\epsilon_{i}^{K_{1}}\leq\epsilon_{i}^{s}$ and $Y_{i}^{K_{1}}\subset Y_{i}$. Hence, it can be obtained that $\mathcal{F}^{K_{1}}=\bigcap^{m}_{i=1}\{x\in X|g_{i}(x,y_{i})\leq-\epsilon^{K_{1}}_{i},\ \forall y_{i}\in Y_{i}^{K_{1}}\}\supset\hat{\mathcal{F}}\neq\emptyset$. Therefore, after at most $K_{1}$ executions of Case I in Algorithm 2, the ($\rm ADRCP^{k}$}) becomes feasible.
2) Exclusion of infinite Case II: Referring to the proof idea of Lemma 2.4 in [29], we exclude the infinite Case II for any agent in the loop of Algorithm 2. For any $i\in\mathcal{V}$, we first consider a sequence of solutions of the ($\rm ADRCP^{k}$}) for this agent. Since $X$ is a compact set, we can select a converging subsequence $x^{k}_{i}\rightarrow\widetilde{x}_{i}$ for agent $i$, where $\widetilde{x}_{i}$ is the convergence point of the subsequence. Assume that at the $k$-th iteration, the point $x^{k}_{i}$ does not satisfy its local constraints $\mathcal{F}_{i}$, and $y^{\max,k}_{i}$ is obtained through solving the (${\rm LLP}_{i}$}). Thus, the following relation holds:
$g_{i}(x^{k}_{i},y^{\max,k}_{i})>0.$
In step 10 of Algorithm 2, the point $y^{\max,k}_{i}$ is added into the finite set of agent $i$. Hence, for any $l>k\geq 0$, it follows that $y^{\max,k}_{i}\in Y^{l}_{i}$. In Section 3.1, the DPG algorithm guarantees the local feasibility of the solution due to the adoption of the Euclidean projection. Hence, for any $l>k\geq 0$, it holds that
$g_{i}(x^{l}_{i},y^{\max,k}_{i})\leq-\epsilon^{l}_{i}.$
$\exists\delta>0\ \ g_{i}(x,y^{\max,k}_{i})\leq-\epsilon^{l}_{i}/2<0\ \ \|x-x^{l}_{i}\|<\delta\ \ \forall l>k\geq 0.$ (16)
As a result of $x^{k}_{i}\rightarrow\widetilde{x}_{i}$, for any $\delta>0$ there is
$\exists K_{2}:\ \|x^{l}_{i}-x^{k}_{i}\|<\delta\ \ \forall l,k:\ l>k\geq K_{2}.$ (17)
Combine the results (16) and (17), we can obtain that
$\exists K_{2}:\ g_{i}(x^{k}_{i},y^{\max,k}_{i})\leq-\epsilon_{i}^{l}/2<0\ \ \forall k:\ k\geq K_{2}.$
Therefore, we can conclude that for any agent $i\in\mathcal{V}$, after at most $K_{2}$ iterations of Algorithm 2, the point $x^{K_{2}}_{i}$ obtained by solving the (${\rm ADRCP}^{K_{2}}$}) lies within its local constraints set $\mathcal{F}_{i}$ in the $({\rm DRCP})$.
3) Exclusion of infinite Case III: Suppose that the (${\rm ADRCP}^{k}$}) is feasible at the $k$-th iteration of Algorithm 2. By using Algorithm 1, each agent can obtain an approximate optimal solution $x^{k}_{i}$ of the (${\rm ADRCP}^{k}$) (cf. Corollary 3.4). Moreover, for any agent $i\in\mathcal{V}$, we have
$\|x^{k}_{i}-x^{k,*}\|=\delta^{k}_{i}$ (18)
where $k\geq 0$. $\delta^{k}_{i}$ is a positive scalar satisfying $0\leq\delta^{k}_{i}\ll\min\left\{\epsilon_{4},\epsilon_{5},\epsilon_{6}\right\}$ as Algorithm 1 is fully iterated by setting small parameters $\epsilon_{1},\epsilon_{2},\epsilon_{3}$, where $\max\left\{\epsilon_{1},\epsilon_{2},\epsilon_{3}\right\}\ll\min\left\{\epsilon_{4},\epsilon_{5},\epsilon_{6}\right\}$. Then, we have $x^{k}_{i}=x^{k,*}+\delta^{k}_{i}v^{k}_{i}$, where $v^{k}_{i}=({x^{k}_{i}-x^{k,*}})/\|x^{k}_{i}-x^{k,*}\|$. For any $i\neq j\in\mathcal{V}$, suppose that they update $z_{i}^{k_{i}+1},z_{j}^{k_{j}+1}$ by Step 13 at the $k_{i}$-th and $k_{j}$-th iterations of Algorithm 2, respectively. Thus, we have
$\hskip-8.53581pt\begin{aligned} \|z_{i}^{k_{i}+1}\!-\!z_{j}^{k_{j}+1}\|&=\!\|x_{i}^{k_{i}}\!-\!x_{j}^{k_{j}}\|\\ &\leq\!\|x^{k_{i},*}\!-\!x^{k_{j},*}\|\!+\!\|\delta^{k_{i}}_{i}v^{k_{i}}_{i}\!-\!\delta^{k_{j}}_{j}v^{k_{j}}_{j}\|\\ &\leq\!\|x^{k_{i},*}\!-\!x^{k_{j},*}\|\!+\!\delta^{k_{i}}_{i}\!+\!\delta^{k_{j}}_{j},\\ &\leq\!\|x^{k_{i},*}\!-\!x^{*}\|\!+\!\|x^{k_{j},*}\!-\!x^{*}\|\!+\!\delta^{k_{i}}_{i}\!+\!\delta^{k_{j}}_{j}.\end{aligned}$ (19)
where the last three inequalities hold based on the triangle inequality. Next, we prove the convergence of the sequence $\{x^{k,*}\}_{k}$. We first make the following settings:
@@
where $\epsilon^{k}_{i}$ and $Y_{i}^{k}$ are the parameters corresponding to the ($\rm ADRCP^{k}$}) at the $k$-th iteration of Algorithm 2. Taking $F(x)$ as the objective function and each of the four sets $\mathcal{F}$, $\mathcal{F}^{k}$, $\mathcal{A}^{k}$ and $\mathcal{B}^{k}$ as the constraint, we obtain the corresponding optimal solutions $x^{*}$, $x^{k,*}$, $x^{k,a}$, $x^{k,b}$ by solving them in turn. Since the relation $Y^{k}_{i}\subset Y^{k+1}_{i}$ holds for any $k\geq 0$ and $i\in\mathcal{V}$, $X$ and $Y_{i}$ are compact, and $g_{i}$ are continuous on $X\times Y_{i}$ for any $i\in\mathcal{V}$, by Lemma 6.1 in [46], it holds that
@@
In addition, since $\epsilon^{k}_{i}\rightarrow 0$ as $k\rightarrow\infty$ for all agents, it follows that $\lim_{k\rightarrow\infty}\mathcal{A}^{k}=\mathcal{F}$. Based on this result, we obtain that